OpenStudy (anonymous):

Center: (16, -9); Radius: 4 Center: (-16, 9); Radius: 4 Center: (16, -9); Radius: 2 Center: (-16, 9); Radius: 2

OpenStudy (anonymous):

Circle S has an equation of (x + 16)2 + (y – 9)2 = 4. What is the center and radius of circle S? Opps heres the question

OpenStudy (anonymous):

option c?

OpenStudy (savvy):

center: (-16,9) radius = 2

OpenStudy (anonymous):

sorry has to be option d

OpenStudy (anonymous):

Ok thanks give each other medals and ill give it to whoever got it right first

OpenStudy (anonymous):

To find the radius in the equations just square the last number... for instance if it is (x + 16)2 + (y – 9)2 = 4 then \[\sqrt{4} = 2\]